OpenStudy (roxy.girl):

Help me solve 6 = -1z + -4

OpenStudy (roxy.girl):

@Qwertty123

rishavraj (rishavraj):

so u need to calculate the calculate the value of z...... fo tht u need to isolate z on one side... first step u can take is add 4 on both sides

OpenStudy (qwertty123):

Hola hola! Reorder the terms: 6 = -4 + -1z Solving 6 = -4 + -1z Move all terms containing z to the left, all other terms to the right.

OpenStudy (qwertty123):

Then Add 'z' to each side of the equation. 6 + z = -4 + -1z + z Combine like terms: -1z + z = 0 6 + z = -4 + 0 6 + z = -4 Add '-6' to each side of the equation. 6 + -6 + z = -4 + -6

OpenStudy (qwertty123):

Next Combine like terms: 6 + -6 = 0 0 + z = -4 + -6 z = -4 + -6 Combine like terms: -4 + -6 =?

OpenStudy (roxy.girl):

-10?

rishavraj (rishavraj):

or u can reach there like add 4 on both sides... u will have 10 = -z and after tht multiply by -1 on bth sides so z = -10

OpenStudy (qwertty123):

yes z = -10! Very well done.

OpenStudy (roxy.girl):

Thanks both of you! I will need help with some more